OpenStudy (cookie_2046):

Solve the equation. If necessary, round your answer to the nearest tenth. 4x^2 = 121 A. 2.75, –2.75 B. 5.5, –5.5 C. 11, –11 D. 30.25, –30.25

OpenStudy (freckles):

do you know how to solve something like x^2=9?

OpenStudy (cookie_2046):

It's D

OpenStudy (freckles):

\[x^2=9 \text{ implies } x=\sqrt{9} \text{ or } x=-\sqrt{9} \text{ are solutions }\]you can simplify those solutions \[x^2=\frac{121}{4} \text{ is what you want \to solve }\] try taking square roots of both sides like I did \[x=\frac{\sqrt{121}}{\sqrt{4}} \text{ or } x=-\frac{\sqrt{121}}{\sqrt{4}}\] do you know the sqrt(121) and the sqrt(4)?

OpenStudy (freckles):

@Cookie_2046 it isn't D

OpenStudy (cookie_2046):

C.

OpenStudy (freckles):

can you tell me what sqrt(121) and sqrt(4) equal?

OpenStudy (cookie_2046):

Yeah.. i don't know that stuff, i'm bad at this

OpenStudy (freckles):

so do you know 11*11=121 and 2*2=4?

OpenStudy (cookie_2046):

Yes

OpenStudy (freckles):

so you should know sqrt(121)=11 and sqrt(4)=2

OpenStudy (freckles):

\[x=\frac{\sqrt{121}}{\sqrt{4}} \text{ or } x=-\frac{\sqrt{121}}{\sqrt{4}}\] simplify further

OpenStudy (cookie_2046):

the second one

OpenStudy (freckles):

no those are both the solutions

OpenStudy (cookie_2046):

Like i said i don't know this stuff what do i simplify?

OpenStudy (freckles):

I was asking you to simplify both using that sqrt(121)=11 and sqrt(4)=2

OpenStudy (cookie_2046):

so 121 x 4?

OpenStudy (freckles):

do you see the division bar (fraction bar) between the sqrt(121) and the sqrt(4)

OpenStudy (freckles):

that doesn't mean multiply

OpenStudy (freckles):

and also what happen to the sqrt( ) part

OpenStudy (freckles):

for example \[x^2=\frac{9}{4} \text{ \implies } x=\frac{\sqrt{9}}{\sqrt{4}} \text{ or } x=-\frac{\sqrt{9}}{\sqrt{4}} \\ \text{ simplifying both of these solutions we get } x=\frac{3}{2} \text{ or } x=- \frac{3}{2}\] see all I did was replace sqrt(9) with 3 and sqrt(4) with 2